An airplane flying parallel to the ground undergoes two consecutive displacements. The first is 79 km at 22.1 west of north, and the second is 108 km at 60.4 east of north.

At what angle east of north is the plane’s total
displacement?
Answer in units of degrees.
Answer #1
x = -79sin(22.1) + 108sin(60.4) = 64.2 km
y = 79cos(22.1) + 108cos(60.4) = 126.5 km
angle = arctan(x/y) = 26.9 degrees